OpenStudy (anonymous):

Average movie prices in the united states are, in general, lower than in other countries. Ot would cost $78.85 to buy 3 tickets in Japan plus 2 in Switzerland. 3 tickets in Switzerland plus 2 tickets in Japan would cost $74.95. How much does tha verage movie ticket cost in each of these countries?

OpenStudy (anonymous):

Lets setup some equations from our known data...

OpenStudy (anonymous):

78.85 = 3J + 2S 74.95 = 3S + 2J as you can see we have two unknowns with two equations and can solve

OpenStudy (anonymous):

we can use the substitution or the elimination method. I prefer substitution. Which would you like, or can you solve it from here on?

OpenStudy (anonymous):

substitution ism uch easier

OpenStudy (anonymous):

can you solve it yourself now?

OpenStudy (anonymous):

no i need help still

OpenStudy (anonymous):

78.85 = 3J + 2S 74.95 = 3S + 2J Solve one of the equations for a variable. 78.85 = 3J + 2S 78.85 - 3J = 2S (78.85/2) - (3/2)J = S Substitute this in to the other equation. 74.95 = 3((78.85/2) - (3/2)J) + 2J Solve for J. Now that we know J, plug that into one of the original equations and solve for S.

OpenStudy (anonymous):

Because both unknowns have a coefficient in front, substitution is actually kind of messy and I would probably go with the elimination method. i will show the steps for this too

OpenStudy (anonymous):

You need to create a common coefficient so one of the terms drops out. (78.85 = 3J + 2S)3 (74.95 = 3S + 2J)2 Now we can solve each equation for the common term 3*78.85 - 9j = 6s 2*74.95 - 4j = 6s Because they equal the same thing we can set them equal to one another 3*78.85 - 9j = 2*74.95 - 4j Solve for j and substitute that back into one of the original equations to get s